OpenStudy (anonymous):

a recipe calls for 2/5 cup +1/4 =

OpenStudy (saifoo.khan):

take the LCM.

OpenStudy (anonymous):

20

OpenStudy (saifoo.khan):

Right^

OpenStudy (anonymous):

2*4=8 1*5=5 8/20+5/20=12/20

OpenStudy (saifoo.khan):

8+5 = 13.

OpenStudy (anonymous):

right opps lol

OpenStudy (anonymous):

13/20

OpenStudy (anonymous):

yups 13/20